JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD

JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD MCA 2022 PROBABILITY AND STATISTICS Previous Question Paper (October/November - 2022 - R13)

MCA | 2022 | I | R13 | October/November - 2022

Answer PaperPaper Code: 811ADMaximum Marks: 60Time: 3 Hours

Answer any five questions. All questions carry equal marks.

Main

112 marks
a

Companies B1, B2 and B3 produce 30%, 45% and 25% of the cars respectively. It is known that 2%, 3% and 2% of these cars produced from B1, B2 and B3 are defective. If a car purchased is found to be defective, what is the probability that this is purchased from company B2?

6 marks

Answer

Let E be the event that a car is defective. Let B1, B2, and B3 be the events that a car is produced by companies B1, B2, and B3 respectively. The given probabilities are: P(B1) = 0.30, P(B2) = 0.45, P(B3) = 0.25. The conditional probabilities of a defective car are: P(E|B1) = 0.02, P(E|B2) = 0.03, P(E|B3) = 0.02.

Using the law of total probability, the total probability of a defective car is:

P(E) = P(B1)P(E|B1) + P(B2)P(E|B2) + P(B3)P(E|B3)
P(E) = (0.30)(0.02) + (0.45)(0.03) + (0.25)(0.02) = 0.006 + 0.0135 + 0.005 = 0.0245

By Bayes' theorem, the probability that the defective car was produced by B2 is:

P(B2|E) = P(B2)P(E|B2) / P(E)
P(B2|E) = 0.0135 / 0.0245 = 135 / 245 = 27 / 49 ≈ 0.551
b

For any three events A, B and C defined on the sample space S verify that

P(A∪B|C) = P(A|C) + P(B|C) - P(A∩B|C)

6 marks

Answer

By the definition of conditional probability, for any event X, P(X|C) = P(X ∩ C) / P(C), provided P(C) > 0. Applying this to the expression P(A ∪ B | C):

P(A ∪ B | C) = P((A ∪ B) ∩ C) / P(C)

Using the distributive law of sets, (A ∪ B) ∩ C = (A ∩ C) ∪ (B ∩ C). Thus:

P(A ∪ B | C) = P((A ∩ C) ∪ (B ∩ C)) / P(C)

Applying the addition theorem for probability, P(X ∪ Y) = P(X) + P(Y) - P(X ∩ Y), where X = A ∩ C and Y = B ∩ C:

P(A ∪ B | C) = [P(A ∩ C) + P(B ∩ C) - P((A ∩ C) ∩ (B ∩ C))] / P(C)
P(A ∪ B | C) = [P(A ∩ C) + P(B ∩ C) - P(A ∩ B ∩ C)] / P(C)

Distributing the denominator P(C):

P(A ∪ B | C) = P(A ∩ C)/P(C) + P(B ∩ C)/P(C) - P(A ∩ B ∩ C)/P(C)

By the definition of conditional probability, this simplifies to:

P(A ∪ B | C) = P(A|C) + P(B|C) - P(A ∩ B | C)
212 marks
a

State and Prove Baye’s Theorem.

6 marks

Answer

Sign in to view the complete answer paper.

b

If A and B are independent events of a sample space S, then show that (i) A and B are independent (ii) A and B are independent (iii) A and B are independent.

6 marks

Answer

Sign in to view the complete answer paper.

312 marks
a

A random variable X has the following probability function:

x01234567
P(X=x)0k2k2k3kk22k27k2+k

Find (i) mean of X. (ii) variance of X.

6 marks

Answer

Sign in to view the complete answer paper.

b

If two cards are drawn from a pack of 52 cards which are diamonds, using Poisson distribution find the probability of getting two diamonds at least 3 times in 51 consecutive trials of two cards drawing each time.

6 marks

Answer

Sign in to view the complete answer paper.

412 marks
a

Is

f(x) = (1/2)x e-x
, when x ≥ 0 can be regarded as a probability function for a continuous random variable? If so, find Mean and Variance of the random variable.

6 marks

Answer

Sign in to view the complete answer paper.

b

A random sample of size 64 is taken from a normal population with mean 51.4 and S.D 68. What is the Probability that the mean of the sample will (i) fall between 50.5 and 52.3 (ii) be less than 50.6.

6 marks

Answer

Sign in to view the complete answer paper.

5

A finite population consists of the numbers 4, 8, 12, 16, 20, 24.

  • List of all possible samples of size 3 without replacement.
  • Find the mean of the population.
  • Find the standard deviation of the population.
  • Calculate the mean of the sampling distribution of means.
  • Find the standard deviation of sampling distributions of means.
12 marks

Answer

Sign in to view the complete answer paper.

6

A professor’s feeling about the mean mark in the final examination in probability of a large group of students expressed subjectively by normal distribution with μ0 = 67.2 and σ0 = 1.5.

  • If the mean mark lies in the interval (65, 70) determine the prior probability the professor should assign to the mean mark.
  • Find the posterior mean μ1 and standard deviation σ1 if the examinations are conducted on a random sample of 40 students yielding mean 74.9 and S.D 7.4.
  • Determine the posterior probability which he will thus assign to mean mark being in the interval (65, 70) using results obtained in (b).
  • Construct a 95% Bayesian interval for μ.
12 marks

Answer

Sign in to view the complete answer paper.

7

Two horses A and B were tested according to the time to run a particular track with the following results:

Horse A: 28, 30, 32, 33, 33, 29, 34

Horse B: 29, 30, 30, 24, 27, 29

Test whether the two horses have the same running capacity.

12 marks

Answer

Sign in to view the complete answer paper.

8

A sample of 12 fathers and their eldest sons gave the following data about their height in inches. Calculate the Coefficient of the rank correlation.

Fathers656367646862706668676971
Sons686668656966686571676870
12 marks

Answer

Sign in to view the complete answer paper.

About this paper

This page contains the solved question paper with answers of PROBABILITY AND STATISTICS from JAWAHARLAL NEHRU TECHNOLOGICAL UNIVERSITY HYDERABAD for MCA 2022 I, October/November - 2022, R13.

For the question paper without answers, switch to the Question Paper tab.

Other papers in PROBABILITY AND STATISTICS

Loading Question Papers

Fetching data from database...